OpenStudy (kittiwitti1):

L̶o̶g̶a̶r̶i̶t̶h̶m̶i̶c̶ ̶D̶i̶f̶f̶e̶r̶e̶n̶t̶i̶a̶t̶i̶o̶n̶ Derivatives of Trigonometric Functions Differentiate the following functions. LaTeX below.

OpenStudy (kittiwitti1):

\[G(x)=\frac{e^{x}-\tan{x}}{\sin{x}+3}\]

OpenStudy (kittiwitti1):

\[G'(x)=\frac{(e^{x}-\tan{x})'(\sin{x}+3)-(e^{x}-\tan{x})(\sin{x}+3)'}{(\sin{x}+3)^{2}}\]

OpenStudy (kittiwitti1):

\[G'(x)=\frac{(e^{x}-\sec^{2}{x})(\sin{x}+3)-(e^{x}-\tan{x})(\cos{x})}{(\sin{x}+3)^{2}}\]Should I keep going or is this good?

zepdrix (zepdrix):

Mmm looks good so far :)

zepdrix (zepdrix):

Are you asking whether or not you should simplify further?

OpenStudy (kittiwitti1):

Yes @zepdrix

zepdrix (zepdrix):

Looks good where it is. Just depends on what your teacher expects I guess..

OpenStudy (irishboy123):

that's the quotient rule, not logarithmic diff are you sure you weren't supposed to take logs first?

OpenStudy (kittiwitti1):

Well... is it \(possible\) to simplify further? Oops, sorry, I meant "Derivatives of Trigonometric Functions," my bad. @IrishBoy123

OpenStudy (coolor):

it doesn't really simplify anymore tho..

zepdrix (zepdrix):

You could expand out all the brackets and some of the trig stuff would change .. I don't think there will be any like-terms to combine though. So it would end up looking "different", not necessarily "simpler" though.

OpenStudy (kittiwitti1):

Alright, then. That works. Thank you @zepdrix :-)

zepdrix (zepdrix):

np kittimittens D: buhaha

OpenStudy (kittiwitti1):

I am not a mitten D:< ! lol

zepdrix (zepdrix):

i call my kitty mr mittens c: had to try it on you lol

zepdrix (zepdrix):

He thinks he's so fancy in his furry tuxedo c:

zepdrix (zepdrix):

lol :3

OpenStudy (kittiwitti1):

Send pix pls

OpenStudy (kittiwitti1):

Lol

zepdrix (zepdrix):

@kittiwitti1 This isn't mr mittens, I don't have a picture of him :P This is his fat friend though. I caught a funny picture of him taking a nap the other day lolol

zepdrix (zepdrix):

@kittiwitti1 XD lol

OpenStudy (kittiwitti1):

THAT IS ADORABLE. @zepdrix